OpenStudy (anonymous):

Evaluate the integral. (8/x - 1/3x^5 + 1/3x^7) dx

OpenStudy (anonymous):

\[\int\limits\frac{8}{x}-\frac{1}{3}\int\limits x^5+\frac{1}{3}\int\limits x^7=8Ln|x|-\frac{1}{18}x^6+\frac{1}{24}x^8\]

OpenStudy (anonymous):

okay so I wrote that but it keeps saying its wrong :/